OpenStudy (anonymous):

find the domain of the function f(x)= \[\sqrt{2x-4}-5\]

OpenStudy (anonymous):

@mathstudent55 hey come

OpenStudy (anonymous):

A. x ≥ 2 B. x > 2 C. x ≥ 4 D. x < -2

OpenStudy (mathstudent55):

Whatever is inside the root must be greater than or equal to zero. Solve the inequality below for x: \( 2x - 4 \ge 0 \)

OpenStudy (mathstudent55):

Add 4 to both sides: \(2x \ge 4 \) Divide both sides by 2: \(x \ge 2 \)